OpenStudy (anonymous):

Write the expression as a single logarithm. Identify any properties used. 3 log4 x + 2 log4 x

OpenStudy (anonymous):

@freckles can you help me again

OpenStudy (freckles):

\[\text{ power rule } r \log_a(x)=\log_a(x^r) \\ \text{ product rule } \log_a(cd)=\log_a(c)+\log_a(d)\]

OpenStudy (freckles):

these are the two rules you will need

OpenStudy (freckles):

for the particular expression

OpenStudy (freckles):

so the first rule says you can do something with the 3 in front of log_4 (x) and the 2 in front of log_4(x)

OpenStudy (freckles):

well there is another way to look at the problem you do have like terms

OpenStudy (anonymous):

So log4 x^3+ log4 x^2

OpenStudy (freckles):

cool

OpenStudy (freckles):

now use product rule

OpenStudy (freckles):

\[\log_a(c)+\log_a(d)=\log_a(c \cdot d)\]

OpenStudy (anonymous):

log4(x^3*x^2) ?

OpenStudy (freckles):

cool

OpenStudy (freckles):

\[\log_4(x^3 \cdot x^2) \\ \log_4(x^{3+2})\]

OpenStudy (anonymous):

Thats all I have to do?

OpenStudy (freckles):

well do 3+2 and then you are done

OpenStudy (anonymous):

Ok. Thank you again lol

OpenStudy (freckles):

\[\log_4(x^5)\]

OpenStudy (freckles):

there is another way to have looked at this problem

OpenStudy (freckles):

you do have like terms above 3u+2u=5u so you had \[3\log_4(x)+2\log_4(x)=(3+2)\log_4(x)=5 \log_4(x) \\ \text{ which this should be fine as an answer too } \\ \text{ or using power rule you have } = \log_4(x^5) \\ \text{ either answer should be fine }\]

OpenStudy (freckles):

anyways be sure to mention the rules you used in your answer because i see your teacher has that as instruction in this problem

OpenStudy (anonymous):

Ok